an electrician requires 12 hours to wire a house, the electricians apprentice can wire a house in 16 hours, after working alone on a job for 4 hours the electrician quits and the apprentice completes the task. how long does it take the apprentice to finish wiring the house?
Answer by Boreal(15235)

the electrician can wire 1/12 of a house per hour.
apprentice can do 1/16 of a house per hour.
(4/12)+(x/16)=1; that means that the two together after x hours will do the whole house.
Multiply by 48 to clear fractions
16+3x=48
3x=32
x=32/3 or 10 2/3 hours.
----------------------------------
check this by realizing the first electrician has done 1/3 the house.
The second one has to do 2/3 of the house and we know it takes 16 hours to do the whole house. 16*(2/3)=32/3 hours.